# AI/ML Based Detection and Categorization of Covert Communication in IPv6 Network

### v1

@@ Abstract
The flexibility and complexity of IPv6 extension headers allow attackers to create covert channels or bypass security mechanisms, leading to potential data breaches or system compromises. The mature development of machine learning has become the primary detection technology option used to mitigate covert communication threats. However, the complexity of detecting covert communication, evolving injection techniques, and scarcity of data make building machine-learning models challenging. In previous related research, machine learning has shown good performance in detecting covert communications, but oversimplified attack scenario assumptions cannot represent the complexity of modern covert technologies and make it easier for machine learning models to detect covert communications. To bridge this gap, in this study, we analyzed the packet structure and network traffic behavior of IPv6, used encryption algorithms, and performed covert communication injection without changing network packet behavior to get closer to real attack scenarios. In addition to analyzing and injecting methods for covert communications, this study also uses comprehensive machine learning techniques to train the model proposed in this study to detect threats, including traditional decision trees such as random forests and gradient boosting, as well as complex neural network architectures such as CNNs and LSTMs, to achieve detection accuracy of over 90%. This study details the methods used for dataset augmentation and the comparative performance of the applied models, reinforcing insights into the adaptability and resilience of the machine learning application in IPv6 covert communication. In addition, we also proposed a Generative AI-assisted interpretation concept based on prompt engineering as a preliminary study of the role of Generative AI agents in covert communication.
keywords:
@@
journal: Computers & Security
[…]
[inst4]organization=Laboratory for Advanced Cybersecurity Research, National Security Agency,country=USA
{highlights}
For covert communications in IPv6 networks, this study applies various AI and machine learning technologies (such as random forests, gradient boosting, and neural networks such as CNN and LSTM) to detect and classify covert communications. In addition to effectively detecting covert communications, it can locate injection fields. The best F1 Score for tasks with or without covert communication is 95%, and the best F1 Score for locating injection field tasks is 99%.
This study comprehensively analyzes IPv6 packet structure, network traffic behavior, and related research weaknesses. This study redesigned the data set and used encryption algorithms for covert communication injection to simulate an attack scenario that is closer to reality.
Generative AI makes systematic improvements by iterative reviewing, adjusting, and honing the classifiers until they reach or surpass the accuracy threshold. This process of refining leads to improved detection and classification accuracy in cybersecurity applications by effectively automating model selection and script improvements utilizing the capabilities of GPT-4-turbo.
1 Introduction
@@
In recent years, with the development of computer computing power and massive research investment in machine learning algorithms, machine learning has succeeded in cybersecurity, including the classification of network traffic [1]. In this context, using machine-learning approaches for detecting packets injected with covert communication and identifying the location of the injected headers become the solution of choice. The goal is to mitigate the threat of injected covert communication when using IPv6. However, attack methods are changing rapidly, and effectively training a machine learning model to detect and classify anomalies requires a large amount of high-quality data. In addition, it also relies on a comprehensive understanding of IPv6’s potential covert communication injection threats to reflect real-world issues.
In this study, we used the prisoner problem as a covert communication scenario[2], where Alice wants to send a message to Bob, and the warden, Wendy, tries to detect it. To address the high-quality dataset required for machine learning training, we developed a real-life IPv6 dataset embedded with hidden communications, mimicking the complexity and subtlety of potential real-world attacks. This work involves a detailed examination of the structure of IPv6 packets to ensure that hidden communications are difficult to detect and decrypt. We evaluate the ability of various machine learning models to detect these covert communications, ranging from traditional algorithms such as random forests and gradient boosting to complex neural networks such as convolutional neural networks (CNN) and long short-term memory networks (LSTM). Our analysis aims to provide a reliable assessment of the effectiveness of these techniques in detecting complex hidden communications within IPv6 [3].
@@
A comprehensive understanding of the potential for IPv6 headers to contain covert traffic and the creation of a more realistic IPv6 covert traffic dataset will provide valuable insights into detecting IPv6 covert communication based on machine learning techniques. Evaluating various machine learning algorithms will contribute to identifying appropriate IPv6 covert communication detection algorithms, making it a potential intrusion detection system (IDS) solution.
2 Related Research
Though critical, the exploration of covert communications within the IPv6 protocols remains under-researched compared to other aspects of network security. The existing research focuses extensively on the potential for embedding covert data within various IPv6 header fields, which the protocol’s expanded architecture has significantly facilitated.
The Study by Lucena et al.[4] has been foundational, identifying up to 22 covert channels within IPv6. These channels leverage underutilized fields such as the Flow Label and Payload Length, offering stealthy methods for transmitting hidden information across networks while evading standard detection mechanisms. Another notable development is the pcapStego tool by Zuppelli and Caviglione (2021), designed to inject covert communications directly into IPv6 headers, simulating real-world covert channel utilization for research purposes [5].
Despite these advancements, a shortage arises in the practical application of such studies, particularly highlighted by Dua et al. [6]. Their research employed pcapStego to craft IPv6 packets with embedded covert channels. While their work on a two-layer detection framework combining Deep Neural Networks (DNN) and Support Vector Machines (SVM) represents a significant technical achievement, whether the crafted covert communications packages reflect real-world threats has been questioned. The resulting covert communications are oversimplified and unrealistic, so these features make the covert packet generated by pcapStego easy to detect. For example, Figure 1 shows a covert communication containing the word “Hello” in the Traffic Class header injected with pcapStego. It can be noticed that the sequence number is reset to 1. Although a Wireshark view of the pcap file would show it as the same flow, these are corrupt TCP packets that, if sent, would result in TCP errors. While injecting the covert channel into the IPv6 header, the pcapStego application created fake traffic that mimicked TCP traffic. As was shown in [7], such traffic is easy to identify. This characteristic will become a feature, unrelated to the actual injected traffic, that makes it easier for machine learning algorithms to identify. This fundamentally undermines the practical significance of their findings and suggests an overestimation of detection capabilities in real-world scenarios.
@@ 2 Related Research
3 Methodology
3.1 Threat Modelling
As shown in Figure 2, our threat modeling is motivated by the idea of the Prisoners’ Problem and the application of Subliminal Channels [2]. In this scenario, two parties, Alice and Bob, want to communicate secretly through IPv6 packets so that Wendy, the warden, will not be able to intercept their communication. Manipulation of several IPv6 header fields, including the Flow Label, Payload Length, and the Source/Destination Address Spaces, may enable covert communication [8]. These fields might contain encrypted data or patterns that a novice observer or a simple detection system would not immediately recognize as anomalous.
Our machine learning model’s primary goal is to enable Wendy to recognize these secret communications and, if feasible, retrieve the hidden information. The model applies sophisticated machine learning techniques as described in [9] to identify discrepancies that indicate covert activities by leveraging patterns found from the analysis of these header fields.
@@ 3.1 Threat Modelling
3.2 Operational Pipeline
Our research employs a sophisticated machine learning pipeline to identify and categorize covert communications within IPv6 network traffic, illustrated in Figure 3. This machine learning pipeline integrates data collection, encryption, preprocessing techniques, and advanced neural network models. Below, we detail each step involved in the pipeline.
3.2.1 Data Collection and Generation
@@
Category Details Count
Training Packets Normal 411,720
@@
Breakdown of Covert Packets HopLimit Encoding 116,628
Address Space 76,575
Length Encrypt 75,823
Flow Label 44,712
@@
Table 1: Summary of Training Packets in the Dataset
  • 1.
  • 2.
    Covert Traffic Creation: To simulate the anomalous or covert class, covert communication traffic was synthetically created. The covert traffic was created using Python scripts that applied different encryption techniques to IPv6 fields. This ensured that the traffic looked authentic. This methodology tests and challenges our machine learning models’ detection capabilities [12].
The dataset used in our machine learning research is summarized in the following table.
@@ 3.2.2 Data Processing
3.2.3 Covert Communication Detection
  • 1.
  • 2.
3.2.4 Covert Communication Classification
  • 1.
3.2.5 Result
  • 1.
  • 2.
  • 3.
4 IPv6 Vulnerability Assessment
4.1 Hop Limit
4.2 Traffic Class
@@ 4.3.1 Pattern Analysis
4.3.2 Special Case with ICMPv6
ICMPv6 packets, particularly those used for “Ping Requests” and “Ping Replies,” exhibit little variation in FlowLabel values, similar to TCP. However, ICMPv6 packets that transmit error messages often have randomly assigned FlowLabel values. This randomness makes them another viable channel for embedding covert messages in a way that is less likely to be detected.
4.3.3 FlowLabel Encryption Strategy
@@
Example of Encrypted FlowLabels: The list in Table 2 provides an example of how encrypted FlowLabels might appear in a stream of covert communications.
@@ 15 FCAC6
Table 2: Example of encrypted FlowLabel values in covert communication packets.
A portion of the hidden message corresponds to each encrypted FlowLabel. The sequence number and the strength of the RC4 encryption ensure that even if the packets are sent in a different order, the recipient will still be able to understand and interpret the message. For instance, “This is a covert communication” is translated by the encrypted FlowLabel values in Table 2.
4.4 Length Field
Since the payload length is specified by the IPv6 Length field—which is not used in checksum calculations—it offers a chance to embed covert communications. This property makes it possible to make changes that are difficult to find using common integrity checks.
4.4.1 Encryption Strategy
@@ 4.5 Address Space
4.5.1 Encryption Methodology
@@ 5.1 Dataset Creation
5.1.1 Covert Packet Integration
  • 1.
@@ 5.1.1 Covert Packet Integration
5.1.2 Covert Communication Challenges
@@
5.2 Machine Learning Models
@@
  • 2.
    Gradient Boosting and XGBoost: By concentrating on fixing mistakes from earlier trees, both models seek to improve predictions progressively. XGBoost, specifically, has been optimized to minimize logarithmic loss across classes [20].
  • 3.
@@
  • 1.
    Deep Neural Network (DNN): Comprises four main layers — 1 dense layer of 256 units, followed by dropout, a second dense layer of 128 units, another dropout, and further dense layers reducing to 2 units for output, designed to extract deep features from complex data patterns [22].
  • 2.
    Convolutional Neural Network (CNN): Includes two convolutional layers and pooling layers, followed by dropout layers and a flatten layer, culminating in dense layers for classification — particularly effective in pattern recognition tasks [22].
  • 3.
    Recurrent Neural Network (RNN) and Long Short-Term Memory (LSTM): The RNN uses two SimpleRNN layers, each followed by dropout, while the LSTM incorporates a bidirectional layer post several dense and dropout layers, enhancing the model’s ability to remember information over extended sequences [23].
  • 4.
@@ 5.3.3 $F_{1}$ Score
[…]
5.3.4 Accuracy

### v2

@@ Abstract
The flexibility and complexity of IPv6 extension headers allow attackers to create covert channels or bypass security mechanisms, leading to potential data breaches or system compromises. The mature development of machine learning has become the primary detection technology option used to mitigate covert communication threats. However, the complexity of detecting covert communication, evolving injection techniques, and scarcity of data make building machine-learning models challenging. In previous related research, machine learning has shown good performance in detecting covert communications, but oversimplified attack scenario assumptions cannot represent the complexity of modern covert technologies and make it easier for machine learning models to detect covert communications. To bridge this gap, in this study, we analyzed the packet structure and network traffic behavior of IPv6, used encryption algorithms, and performed covert communication injection without changing network packet behavior to get closer to real attack scenarios. In addition to analyzing and injecting methods for covert communications, this study also uses comprehensive machine learning techniques to train the model proposed in this study to detect threats, including traditional decision trees such as random forests and gradient boosting, as well as complex neural network architectures such as CNNs and LSTMs, to achieve detection accuracy of over 90%. This study details the methods used for dataset augmentation and the comparative performance of the applied models, reinforcing insights into the adaptability and resilience of the machine learning application in IPv6 covert communication. We further introduce a Generative AI-driven script refinement framework, leveraging prompt engineering as a preliminary exploration of how generative agents can assist in covert communication detection and model enhancement.
Index Terms:
@@
I Introduction
@@
In recent years, advances in computing power and significant research investment have propelled machine learning to the forefront of cybersecurity, particularly in the classification and analysis of network traffic [1]. As a result, machine learning techniques have become the preferred approach for detecting packets containing covert communications and pinpointing the specific header fields used for such injections in IPv6 traffic. The primary objective is to mitigate the risks posed by covert communication attacks within IPv6-enabled networks. However, with the rapid evolution of attack techniques, effectively training machine learning models for this purpose demands access to large volumes of high-quality data, as well as a comprehensive understanding of the diverse methods by which covert communications can be embedded within IPv6, in order to realistically capture and address emerging threats.
In this study, we adopt the prisoner’s problem framework [2] to model a covert communication scenario, where Alice attempts to transmit a hidden message to Bob while an adversary, Wendy, monitors the traffic for evidence of covert activity. To support machine learning-based detection, we construct a realistic IPv6 dataset with embedded covert communications. The dataset is designed to reflect practical usage patterns and structural variability in IPv6 header fields, thereby capturing the complexity and subtlety of real-world covert channels.
Using this dataset, we evaluate a range of machine learning models, including random forests, gradient boosting, convolutional neural networks (CNN), and long short-term memory networks (LSTM). These models are assessed for their ability to detect covert communications embedded within IPv6 packets. The experimental results offer insights into the relative effectiveness of different algorithms and inform the selection of models suitable for IPv6 traffic analysis. This work contributes to the development of machine learning-based intrusion detection systems (IDS) [3, 4] by providing both a representative data set and an empirical benchmark for covert channel detection in IPv6 environments.
II Related Works
While the security of IPv6 has been extensively studied in various domains, the specific issue of covert communications within the protocol remains comparatively underexplored. Prior research has primarily concentrated on the feasibility of embedding hidden data in IPv6 header fields, a possibility enabled by the protocol’s expanded header structure and increased field flexibility.
A recent survey by Khadse and Dakhane consolidates 2015–2024 work on covert-channel construction and detection, emphasizing bandwidth–undetectability trade-offs and persistent dataset scarcity [5]. The Study by Lucena et al. [6] has been foundational, identifying up to 22 covert channels within IPv6. These channels leverage underutilized fields such as the Flow Label and Payload Length, offering stealthy methods for transmitting hidden information across networks while evading standard detection mechanisms. Another notable development is the pcapStego tool by Zuppelli and Caviglione (2021), designed to inject covert communications directly into IPv6 headers, simulating real-world covert channel utilization for research purposes [7]. Another recent advance is the open-source IPv6CC suite, which provides penetration testers and researchers with multiple methods to inject and evaluate covert channels in IPv6 networks using various header fields and injection strategies [8]. IPv6CC enables assessment of organizational security posture against stegomalware and exfiltration threats, and allows for practical benchmarking of detection systems, including their ability to withstand both naive and sophisticated attacks. Suprun et al. propose a Flow Label centric steganographic model secured with ECDH key exchange and ECDSA signatures, reducing encode/decode latency versus earlier CBC-RC6 variants while retaining a 20-bit-per-packet capacity [9]. On the detection side, bccstego introduces a scalable eBPF-based framework that inspects IPv6 header fields in real time to collect statistical indicators and reveal covert channel activity [10]. By leveraging efficient binning and a low processing footprint, bccstego can be deployed in production environments for proactive detection without adversely impacting legitimate traffic or system resources.
@@ II Related Works
Despite these advancements, Dua et al. [11] reveal a gap in practical applicability, as their study is based on idealized assumptions that limit its relevance to real-world deployment scenarios. Their research employed pcapStego to craft IPv6 packets with embedded covert channels. While their work on a two-layer detection framework combining Deep Neural Networks (DNN) and Support Vector Machines (SVM) represents a significant technical achievement, whether the crafted covert communications packages reflect real-world threats has been questioned. The resulting covert communications are oversimplified and unrealistic, so these features make the covert packet generated by pcapStego easy to detect. For instance, Figure 1 illustrates a covert communication scenario in which the word “Hello” is embedded within the IPv6 Traffic Class field using the pcapStego tool. Although Wireshark displays these packets as part of a continuous TCP flow, the sequence number has been manually reset to 1. This inconsistency renders the packets invalid from the perspective of the TCP protocol stack and would likely result in connection errors if transmitted over a real network. While injecting the covert channel into the IPv6 header, the pcapStego application created fake traffic that mimicked TCP traffic. As discussed in [12], this type of traffic is relatively easy to identify. The resulting artifacts, which are unrelated to the content of the injected data, can inadvertently serve as distinguishing features for machine learning algorithms. This reliance on such superficial characteristics undermines the practical relevance of the findings and may lead to an overestimation of detection performance in real-world scenarios.
@@
Recent field measurements by Mazurcyk et al. demonstrate that multiple covert techniques are already deployed in operational networks and benchmark detection performance across six widely used hiding tools. These real-world observations underscore the need for datasets and detection methods that reflect live traffic conditions [13]. Additionally, Dua et al. (2023) introduced the DICCh-D framework, using deep neural networks on blended CAIDA normal traffic and realistic covert samples, to improve detection accuracy to over 99.5% [14]. Recent work by Bedi et al. [15] introduced SPYIPv6, a two-layer framework that uses KNN-based classification to not only detect covert IPv6 packets but also accurately localize the header fields (or combinations thereof) used for data hiding. SPYIPv6 demonstrates high accuracy (99.85%) and fast detection on blended CAIDA and synthetically generated covert traffic, representing a significant advancement in practical detection and localization of multi-field IPv6 covert channels.
A complementary rule-based approach, CC-Guard [16], matches IPv6 header values against learned normal-traffic templates and achieves real-time detection of ten storage channels with negligible runtime overhead. Further advancing the field, Zhang et al. [17] introduce a self‑attention mechanism‑based model for detecting multi‑field covert channels in IPv6, covering both base and extension headers. Their model employs multi‑head attention to correlate sub‑field embeddings and accurately locate covert data across 23 channel types. Evaluated on a publicly available IPv6 covert channel dataset, it demonstrated precision of 97. 13% with a false positive rate of just 6. 3%, significantly outperforming previous approaches such as BNSCNN and DICChD in both detection breadth and localization accuracy. These studies reinforce the importance of quality and representativeness in covert traffic datasets.
III Methodology
III-A Threat Modelling
@@
As shown in Figure 2, our threat modeling is motivated by the idea of the Prisoners’ Problem and the application of Subliminal Channels [2]. In this scenario, two parties, Alice and Bob, want to communicate secretly through IPv6 packets so that Wendy, the warden, will not be able to intercept their communication. Manipulation of several IPv6 header fields, including the Flow Label, Payload Length, and the Source/Destination Address Spaces, may enable covert communication [19, 13, 20, 21]. These fields might contain encrypted data or patterns that a novice observer or a simple detection system would not immediately recognize as anomalous.
Our primary goal is to enable Wendy to recognize these secret communications and, and retrieve the hidden information with machine learning models. These models applies sophisticated machine learning techniques to identify discrepancies that indicate covert activities using patterns found from the analysis of these header fields [22, 23].
@@
III-B Operational Pipeline
Our research employs a comprehensive machine learning pipeline to detect and classify covert communications embedded within IPv6 network traffic, as illustrated in Figure 3. This pipeline encompasses stages such as data collection, encryption, preprocessing, and the application of advanced neural network models. Below, we provide an overview of the five key steps involved in this process. Each step will be discussed in greater detail in the subsequent subsections.
  1. 1.
    Data Collection and Generation: Covert Communication Traffic are Generated and carefully crafted with various encryption methods of different IPv6 fields and python script to automate the process.
  2. 2.
    Data Processing: The vulnerable IPv6 fields are selected for analysis through Feature Selection. Standard data processing steps like Normalization or Label Encoding may or may not be performed, depending on the nature of the IPv6 field.
  3. 3.
    Covert Communication Detection: A Binary Classification Model is trained to classify IPv6 Network Packets based on Normal Behavior. After prediction, the normal and suspected covert communication IPv6 packets are separated, where the Suspected Covert Communication Packets will undergo further analysis.
  4. 4.
    Covert Communication Classification: After separating the covert communication data packets from the normal ones, the covert communication packets are passed to a multiclass classification model. The multi-class classification model further classifies the covert comm packets, based on the IPv6 fields that were exploited for covert communication
  5. 5.
    Result: After the complete cycle, the covert communication and normal data packets in IPv6 are sent to the Generative AI agent where the covert communication packets are analyzed to see whether there are more patterns that can be used to identify and categorize covert communication packets, and based on the newly found products and classification results, the python scripts for detection and classification both are revised and refined and the iteration goes on.
III-B1 Data Collection and Generation
@@
  • Covert Traffic Creation: To simulate the anomalous or covert class, covert communication traffic was created synthetically. The covert traffic was created using Python scripts that applied different encryption techniques to IPv6 fields. This ensured that the traffic looked authentic. This methodology tests and challenges our machine learning models’ detection capabilities.
Category Count
Normal 411,720
@@
   – HopLimit Encoding 116,628
   – Address Space 76,575
   – Length Encrypt 75,823
   – Flow Label 44,712
@@
TABLE I: Summary of Training Packets in the Dataset
The dataset used in our machine learning research is summarized in Table I.
@@ III-B2 Data Processing
III-B3 Covert Communication Detection
III-B4 Covert Communication Classification
III-B5 Result
IV IPv6 Vulnerability Assessment
IV-A Hop Limit
@@
IV-B Traffic Class
@@ IV-C1 Pattern Analysis
IV-C2 Special Case with ICMPv6
ICMPv6 packets used for diagnostic purposes, such as Echo Requests and Replies, typically exhibit consistent or default Flow Label values, similar to TCP traffic. In contrast, ICMPv6 error messages often show implementation-dependent Flow Label values, which can sometimes appear random [32]. This inconsistency opens the possibility of using the Flow Label field in such packets as a covert communication channel with a lower risk of detection.
IV-C3 Encryption Strategy on Flow Label
@@
Example of Encrypted Flow Labels: The list in Table II provides an example of how encrypted Flow Labels might appear in a stream of covert communications.
@@ 15 FCAC6
TABLE II: Example of encrypted Flow Label values in covert communication packets.
A portion of the hidden message corresponds to each encrypted Flow Label. The sequence number and the strength of the RC4 encryption ensure that even if the packets are sent in a different order, the recipient will still be able to understand and interpret the message. For instance, “This is a covert communication” is translated by the encrypted Flow Label values in Table II.
IV-D Length Field
Since the payload length is specified by the IPv6 length field, which is not used in checksum calculations, it offers a chance to embed covert communications. This property makes it possible to make changes that are difficult to find using common integrity checks.
IV-D1 Encryption Strategy on Length Field
@@ IV-E Address Space
IV-E1 Encryption Strategy on Address Space
@@ V-A Dataset Creation
V-A1 Covert Packet Integration
@@
V-A2 Covert Communication Challenges
@@
V-B Machine Learning Models
@@
Figure 5: Schematic of the neural network architectures explored: (a) a fully connected DNN, (b) a CNN with 1D convolutions, (c) an RNN based on SimpleRNN cells, a (d) GCN and (e) a bidirectional LSTM network.
@@
  • Gradient Boosting and XGBoost: By concentrating on fixing mistakes from earlier trees, both models seek to improve predictions progressively. XGBoost, specifically, has been optimized to minimize logarithmic loss across classes [34, 35].
@@
  • Deep Neural Network (DNN): Comprises four main layers: 1 dense layer of 256 units, followed by dropout, a second dense layer of 128 units, another dropout, and additional dense layers reducing to 2 units for output, designed to extract deep features from complex data patterns [37].
  • Convolutional Neural Network (CNN): Includes two convolutional layers and pooling layers, followed by dropout layers and a flatten layer, culminating in dense layers for classification, particularly effective in pattern recognition tasks [38].
  • Recurrent Neural Network (RNN) and Long Short-Term Memory (LSTM): The RNN uses two SimpleRNN layers, each followed by dropout, while the LSTM incorporates a bidirectional layer post several dense and dropout layers, enhancing the model’s ability to remember information over extended sequences [39, 40].
@@ V-C3 $F_{1}$ Score
$\text{$F_{1}$ Score}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)
V-C4 Accuracy
@@
V-D Generative AI Assisted Refinement
In this work, we refined Python scripts related to covert communication detection using the GPT-4-turbo model as a generative AI agent expert in code generation [42, 43]. Utilizing generative AI, the classifier models underwent iterative refinement, with a binary classifier for detection and a multi-class classifier for identifying the type of covert communication. In accordance with the approach outlined by [44], the generative AI agents aimed to iteratively refine the Python scripts until their performance surpassed a specified threshold. Figure 6 illustrates the process.
Figure 6: AI Agent-Driven Script Refinement Pipeline
The task assignment includes detailed information about the dataset, such as column properties, specific IPv6 fields vulnerable to covert communication, and other contextual information that can affect the detection process. The user prompt contains the path to the dataset, classification results (provided as data frames), Python scripts that need to be refined, and the classification accuracy threshold. Generative AI makes systematic improvements by iteratively reviewing, adjusting, and optimizing the classifiers until they reach or exceed the accuracy threshold. This process of refinement leads to improved detection and classification accuracy in cybersecurity applications by effectively automating model selection and script improvements using the capabilities of GPT-4-turbo.
[…]
This work focuses on refining an existing Python-based detection pipeline (covert_ipv6.py) for identifying and classifying covert communication within IPv6 traffic. The dataset, located at dataset.csv, includes labeled IPv6 packets with features extracted from various header fields, such as Flow Label, AddressSpacePattern, PayloadLength, and others. The classification results, including binary and multiclass outcomes, are stored in results.txt, while domain-specific knowledge about covert channel strategies is available in knowledge.txt.
The refinement process involves the following objectives:
  1. 1.
    Code Enhancement: Improve the detection and classification logic in covert_ipv6.py, particularly feature engineering, data preprocessing, and classifier selection.
  2. 2.
    Binary Classification: Detect whether a given IPv6 packet is part of a covert channel or not. This involves improving the model’s accuracy, precision, recall, and F1-score.
  3. 3.
    Multiclass Classification: For packets classified as covert, further categorize them into specific covert types based on their manipulation strategies (e.g., Flow Label-based, LengthField-based).
  4. 4.
    Knowledge-Guided Feature Augmentation: Incorporate human knowledge from knowledge.txt to derive new features, rules, or validation criteria. Examples include entropy of field values or irregular sequencing of header bytes.
  5. 5.
    Evaluation and Visualization: Quantitatively evaluate the improvements via standard classification metrics and visualize the performance using confusion matrices and ROC curves.
[…]
VI Results and Discussions
This section presents a comprehensive evaluation of the machine learning models deployed to detect and classify covert communications in IPv6 traffic. The analysis is structured around two primary tasks: binary classification for identifying whether a packet contains covert communication and multiclass classification for determining the specific type of covert channel utilized. Standard performance metrics (including precision, recall, $F_{1}$ score, and precision) are used to assess the effectiveness of each model.
Figures 7 and 8 illustrate comparative performance between models using bar graph visualizations. The analysis encompasses traditional ensemble methods including RF, GB, XGB, and LGB, as well as neural network architectures such as DNN, CNN, RNN, and GCN. The following subsections detail the performance of each model in the two classification tasks, providing insights into their relative strengths, limitations, and suitability for different detection scenarios.
VI-A Binary Classification
Among the evaluated models, Random Forest (RF), Extreme Gradient Boosting (XGB), and Light Gradient Boosting (LGB) showed consistent performance across multiple evaluation metrics. In particular, RF and XGB achieved relatively higher accuracy and $F_{1}$ scores compared to the other methods, indicating their suitability for the binary classification task in this context.